A rigid cylinder containing 0.01 m3 of carbon dioxide at 1 bar and 300 K 

is heated until temperature becomes 400 K. Determine: The heat 

supplied.


1
Expert's answer
2021-06-15T17:00:16-0400

Gives

Volume (V)=0.01"m^3"

Pressure (P)=1bar=100Kpascal

∆T=(T2-T1)=400-300=100K

R=4.1240Kpa m3/Kg-K

"c_v=37.35KJ\/kg-K"

"m=\frac{P_1V}{RT_1}"

Put value


"m=\frac{100kpa \times 10^{-2}m^2}{4.12 \frac{kpa m^3}{kg-K }\times300 K}=8.09\times10^{-4} kg"

heat transfer

"Q=mc_v\u2206T"



"Q=8.09\times10^{-4} \times37.35\times100=3.02J"